# The vacuum stability and the hierarchy problem in a fermionic dark matter model

### v2

@@ 3.1 Relic density
where $H$ is the Hubble parameter and $n_{\psi}^{eq}\sim(m_{\psi}T)^{3/2}e^{-m_{\psi}/T}$ is the particle density before particles get out of equilibrium. In our model, the DM fermions $\psi^{1,2}$ possess identical masses and contribute equally to the relic abundance of fermionic DM ($\psi$). The relevant Feynman diagrams for computation of the DM annihilation cross section is depicted in Fig1.
@@ 3.4 Results
We also consider three coupling values of 0.02, 0.6 and 2 for $g_{v}$. The first constraint for the parameter space under consideration is the value of the relic density of DM, which must be in agreement with the current observed value by the PLANCK collaboration[45]. We calculate the relic density numerically by implementing the model into micrOMEGAs[46]. The second constraint is related to direct detection experiments, where we use the results of the XENONnT experiment[47]. The neutrino floor, which is an irreducible background resulting from the scattering of SM neutrinos by nuclei, is also considered[48]. Finally, we use the ATLAS experimental results that put an upper limit on the invisible Higgs decay to constrain the parameter space[44]. Figure 3(a) shows the parameter space in agreement with the Relic density. Due to the constraint of invisible Higgs decay, the $M_{V}$ must be greater than $26~\rm GeV$. Figure 3(b) also shows the allowable points in agreement with the relic density and direct detection. In both of these diagrams, $g_{v}$ is assumed to be equal to 0.02. In Figure 3(c), the permissible parameter space in agreement with the relic density is plotted, and in Figure 3(d), the acceptable points in agreement with the relic density and direct detection for $g_{v}=0.6$ are plotted. A similar analysis is plotted in Figures 3(e) and 3(f) for $g_{v}=2$. Of course, all diagrams include the constraint of invisible Higgs decay. As can be seen, the model is in agreement with the phenomenological constraints for a wide range of parameter space.
@@ 4 Renormalization group equations and vacuum stability
We choose two benchmark points, which are given in Table 2 and are in agreement with all the experimental constraints presented. Using the obtained RGEs, the evolution of couplings with energy can be obtained. Figures 4 and 5 show the evolution of these couplings. As can be seen, all theoretical constraints in equations 2.16 and 2.17 are verified up to the Planck scale and the stability of the model vacuum is fully guaranteed.
#\# $M_{V}(GeV)$ $M_{\psi}(GeV)$ $M_{H_{2}}(GeV)$ $g_{v}$ $g_{s}$ $\sigma_{DM-N}(cm^{2})$
1 1788 290.3 612.1 0.6 0.13 $3.39\times 10^{-47}$
2 37.39 188.5 406.1 0.02 0.14 $8.05\times 10^{-47}$
Table 2: Two benchmark points with DM parameters.
Figure 4: Running of couplings of the model up to Planck scale for Benchmark point 1 in Table 2.
Figure 5: Running of couplings of the model up to Planck scale for Benchmark point 2 in Table 2.
@@ 5 Veltman condition
#\# $M_{V}(GeV)$ $M_{\psi}(GeV)$ $M_{H_{2}}(GeV)$ $g_{v}$ $\sigma_{DM-N}(cm^{2})$
1 1499 1149 2297 0.5 $5.65\times 10^{-46}$

### v3

@@ 3.1 Relic density
where $H$ is the Hubble parameter and $n_{\psi}^{eq}\sim(m_{\psi}T)^{3/2}e^{-m_{\psi}/T}$ is the particle density before particles get out of equilibrium. In our model, the DM fermions $\psi^{1,2}$ possess identical masses and contribute equally to the relic abundance of fermionic DM ($\psi$). The relevant Feynman diagrams for computation of the DM annihilation cross section is depicted in Fig1. In figure 2 we have plotted DM relic density diagram versus the DM mass. As can be seen there are two minima observed in the DM relic density around 62 GeV and 120 GeV. These reductions in relic density at these points are associated with the resonances of $H_{1}$ and $H_{2}$, which is a general feature of purely scalar-portal DM scenarios that $\Omega h^{2}\sim 0.1$(observed relil density) majorly occurs near the resonance dips.
@@ 3.3 Invisible Higgs decay
We also used constraints related to mixing between scalars ($sin\alpha$) in checking our results. Non-observation of DM at the LHC corresponds to an upper bound on the value of $sin\alpha$. The Higgs boson data from LHC at 7 and 8 TeV sets a constraint of $|sin\alpha|<0.36$ at 95% C.L., independently of the $H_{2}$ mass[45]. There is also a mass-dependent constraint, which requires $sin\alpha<0.32$ for $m_{H2}>200GeV$ and $sin\alpha<0.2$ for $m_{H2}>400GeV$, mostly coming from restrictions on the NLO corrections to the mass of the W boson[46].
@@ 3.4 Results
We also consider three coupling values of 0.02, 0.6 and 2 for $g_{v}$. The first constraint for the parameter space under consideration is the value of the relic density of DM, which must be in agreement with the current observed value by the PLANCK collaboration[47]. The value of $\Omega_{DM}$($\Omega_{\psi}$), as reported by the PLANCK collaboration, carries an estimated theoretical uncertainty around 10 percent, $0.11\leq\Omega_{DM}h^{2}\leq 0.13$. We calculate the relic density numerically by implementing the model into micrOMEGAs[48]. The second constraint is related to direct detection experiments, where we use the results of the XENONnT and LUX-ZEPPELIN(LZ) experiments[49, 50]. The LZ experiment places the strongest possible constraint on direct detection results. The neutrino floor, which is an irreducible background resulting from the scattering of SM neutrinos by nuclei, is also considered[51]. Finally, we use the ATLAS experimental results that put an upper limit on the invisible Higgs decay to constrain the parameter space[44]. Figure 4(a) shows the parameter space in agreement with the Relic density. Due to the constraint of invisible Higgs decay, the $M_{V}$ must be greater than $26~\rm GeV$. Figure 4(b) also shows the allowable points in agreement with the relic density and direct detection. In both of these diagrams, $g_{v}$ is assumed to be equal to 0.02. In Figure 4(c), the permissible parameter space in agreement with the relic density is plotted, and in Figure 4(d), the acceptable points in agreement with the relic density and direct detection for $g_{v}=0.6$ are plotted. A similar analysis is plotted in Figures 4(e) and 4(f) for $g_{v}=2$. Of course, all diagrams include the constraint of invisible Higgs decay. As can be seen, the model is in agreement with the phenomenological constraints for a wide range of parameter space.
@@ 4 Renormalization group equations and vacuum stability
We choose one representative benchmark which is given in Table 2 and is in agreement with all the experimental constraints presented. Using the obtained RGEs, the evolution of couplings with energy can be obtained. Figures 5 show the evolution of these couplings. As can be seen, all theoretical constraints in equations 2.16 and 2.17 are verified up to the Planck scale and the stability of the model vacuum is fully guaranteed.
#\# $M_{V}(GeV)$ $M_{\psi}(GeV)$ $M_{H_{2}}(GeV)$ $g_{v}$ $g_{s}$ $\Omega_{DM}h^{2}$ $\sigma_{DM-N}(cm^{2})$ $\sin\alpha$
1 94.07 381.5 766.9 0.02 0.11 0.114 $9.83\times 10^{-48}$ 0.05
Table 2: The benchmark point with DM parameters.
Figure 5: Running of couplings of the model up to Planck scale for Benchmark point in Table 2.
@@ 5 Veltman condition
Of course, it is important to note that what distinguishes the Veltman’s approach in our new model from the results of previously presented models is the new parameter space obtained by simultaneously examining the three issues of DM, vacuum stability, and the hierarchy problem. In fact, adding new symmetries or fields to the SM leads to different phenomenology for DM, different potentials for the model, or leads to different RGEs, which differentiates its result from other models.
#\# $M_{V}(GeV)$ $M_{\psi}(GeV)$ $M_{H_{2}}(GeV)$ $g_{v}$ $\Omega_{DM}h^{2}$ $\sigma_{DM-N}(cm^{2})$ $\sin\alpha$
1 1507 1145 2325 0.27 0.13 $4.51\times 10^{-47}$ 0.04